OpenStudy (anonymous):

The heat H in watts generated by an electric heating element varies directly as the square of the current i in amperes. A heating element produces thermal energy at a rate of 648 watts when the current is 0.9 amperes. What rate of heating in watts will be generated by a current of 1.4 amperes?

OpenStudy (kropot72):

\[648=0.9^{2}\times R\] where R is a constant \[R=\frac{648}{0.9^{2}}\] When the current is increased to 1.4 amperes the watts produced will be\[\frac{648}{0.9^{2}}\times 1.4^{2}\]

OpenStudy (anonymous):

thatk you

OpenStudy (kropot72):

You're welcome :)